# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

### v2

@@ 6.4 Limitations
One such limitation concerns sales data, as we only collected data from two marketplaces. Tracking every single transfer that is also a sale is burdensome because each marketplace has its own smart contracts and sales protocol. Moreover, these might even change over time. For instance, we noticed a sales drop that lasted several days on OpenSea in May 2022, which upon further inspection resulted from a migration to their current protocol called SeaPort. Nonetheless, the marketplaces we selected should cover the vast majority of wearable sales, as Decentraland’s Marketplace is the official and default venue, and OpenSea is the largest NFT marketplace and was the initial default venue to sell wearables by DG, which now has its own marketplace.
Another potential limitation regards in-world tracking of users. We did not use the term active users in our analysis because we only considered the presence of a given account address in the log events, i.e., we did not filter out potential bots or users that could be deemed inactive. In line with other kinds of social media, bots are also a plague in social virtual worlds, with automated avatars having a noticeable impact on these Varvello & Voelker (2010). In addition, a user might be authentic (i.e, not a bot) but be away from keyboard while their avatar is still present in the virtual world. Determining an avatar who represents an “authentic active user”, however, depends on an arbitrary set of characteristics and mechanisms. Indeed, calculating the number of daily active users on Decentraland has sparked much public debate, as different definitions and mechanisms might result in vastly different numbers Thompson et al. (2022). Our approach, on the other hand, is relatively simplistic, but at the same time it better responds RQ2, as in any case these avatars —active or inactive, human or bot— occupy space and resources on Decentraland’s virtual world.
7 Conclusions

### v3

@@ D.   Limitations
One such limitation concerns sales data, as we only collected data from two marketplaces. Tracking every single transfer that is also a sale is burdensome because each marketplace has its own smart contracts and sales protocol. Moreover, these might even change over time. Nonetheless, the marketplaces we selected should cover the vast majority of wearable sales, as Decentraland’s Marketplace is the official and default venue, and OpenSea is the largest NFT marketplace and was the initial default venue to sell wearables by DG, which now has its own marketplace.
Another potential limitation regards in-world tracking of users. We did not use the term active users in our analysis because we only considered the presence of a given account address in the log events, i.e., we did not filter out potential bots or users that could be deemed inactive. In line with other kinds of social media, bots are also a plague in social virtual worlds, with automated avatars having a noticeable impact on these [52]. In addition, a user might be authentic (i.e, not a bot) but be away from keyboard while their avatar is still present in the virtual world. Determining an avatar who represents an “authentic active user”, however, depends on an arbitrary set of characteristics and mechanisms. Our approach, on the other hand, is relatively simplistic, but at the same time it better answers RQ2, as in any case these avatars —active or inactive, human or bot— occupy space and resources on Decentraland’s virtual world.
Finally, our focus on quantitative methods and the lack of qualitative methods, such as user interviews and questionnaires, could be a limitation for investigating the motivations behind users’ actions. We believe that using a quantitative approach based on public data to identify global patterns of actual user behavior is appropriate to answer our study’s RQs. Nevertheless, future studies could also adopt a more comprehensive and multifaceted approach using mixed methods. For instance, by corroborating data analyses such as ours with surveys to obtain data about the users’ self-perceived behavior, or even with ethnographic analyses [5, 8] to delve into the details of users’ behaviors in social virtual worlds.
VII.   Conclusions
@@
Funding
This research received no specific grant from any funding agency in the public or private sectors.
Authors’ Contributions
All authors have participated in drafting the manuscript. All authors read and approved the final version of the manuscript.
Conflict of Interest
The authors declare no conflict of interest.
Data Availability
The data supporting the findings of this study are available on FigShare at: https://doi.org/10.6084/m9.figshare.28596923.
Ethical Statement
Ethical approval was not required for the present study due to its use of anonymous and public data. In particular, the vast majority of our data was obtained from blockchain platforms with permissive software and data licenses. For these data, there is not personally identifiable information of the analyzed users, as these are anonymously identified via a unique string of hexadecimal characters that was cryptographically generated.
Declaration of AI Usage
No AI tools were used in the creation of this manuscript.